The box is pulled by the force Fx, which causes it to accelerate a = 2 m / s². Find a) the work performed by the force Fx, when the box is moved by 5 m; b) work performed by friction force, if the coefficient of friction is 0.4.


1
Expert's answer
2021-04-08T08:51:21-0400

I assume "F_x" is the net force acted on the body. Then the mass of the box can be found from the second Newton's law:


"m = \dfrac{F_x}{a}"

where "a = 2m\/s^2" is the acceleration of the box.

a) By definition, the work of the constant force applied along the body's path is given as follows:


"A _x= F_xs"

where "s = 5m". Thus, obtain:


"A_x = 5F_x\space\space (J)"

b) By definition, the friction force on the horizontal surface is given as follows:


"F_f = \mu mg"

where "\mu = 0.4" is the coefficient of friction, and "g = 9.8m\/s" is the garaviational acceleration. Thus, the work of the friction force is:


"A_f = \mu mgs = \mu\dfrac{F_x}{a}g\\
A_f = 0.4\cdot \dfrac{F_x}{2}\cdot 9.8 = 1.96F_x \space \space (J)"

Answer. a) 5Fx, b) 1.96Fx.